# : Generalizing Large Language Models for E-commerce

### v1

@@ Appendix G OOD Evaluation
G.1 Results on Unseen Instructions
TableΒ A15 exhibits the results of evaluating on OOD test sets with unseen instruction. As shown in tableΒ A15, models perform better when training on with diverse instructions. fine-tuned on with diverse instruction enhances the generalizability of models to the unseen instruction.
@@ G.1 Results on Unseen Instructions
In this table, ``single" and ``diverse" indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model tuned over single and diverse instructions on each task is in bold.
G.2 Results on Different Base Models
TableΒ A16 shows the OOD evaluation results of with different base models. With high-quality dataset, models achieve good performance under different base models.
@@ G.2 Results on Different Base Models
In this table, ``Base Model" presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
G.3 Results on Models Tuned using All Tasks and Individual Tasks
Demonstrated in TableΒ A17, when generalizing to OOD test, fine-tuned on every single task can better transfer the knowledge to the specific task than fine-tuned on the task combination.
| Model | Training | ||||||
@@ G.3 Results on Models Tuned using All Tasks and Individual Tasks
In this table, ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold.
Appendix H Complete Results for the Analysis on Training Data Size
@@
This section exhibits the comprehensive results for analyzing how the training data size influences the performance of models. With the total 92K samples in , we assess data scaling with sizes of 1K, 10K, and 47K. The 1K and 10K samples are collected by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are assembled by randomly selecting 5K samples from each task except for , for which all its 2K training samples are included. As shown in tableΒ A18 and tableΒ A19, performances exhibit an upward trend along with the increase in data size. Notably, the F1 score of the task experiences an initial drop followed by a subsequent rise with the data size increasing. This behavior could be attributed to the imbalanced data, making the models tend to randomly guess when the data size is small. As the data size increases, the models would predict the dominant label, leading to a temporarily low F1 score. With continued data growth, the models acquire sufficient knowledge to generate accurate predictions. In general, the large-scale training data plays a crucial role in developing effective e-commerce LLMs, underscoring the significance of our extensive, comprehensive, and high-quality e-commerce instruction dataset .
Appendix I Case Study
We show the predictions from and the best-performing general-purpose LLM GPT-4 Turbo on each of the 10 tasks. From the following examples, we note that GPT-4 Turbo struggles with performing e-commerce tasks. For instance, GPT-4 Turbo shows difficulties in differentiating products (), estimating users' intent (), and predicting the relevance between user queries and products (). We also observe that by instruction tuning, captures the knowledge required and could perform e-commerce tasks effectively.
I.1 Attribute Value Extraction ()
@@ I.2 Product Matching ()
Product 2: title - omni web 5.0, description - sure you can use a standard web browser with standard features. but you didn't choose a standard software experience - you chose the mac. why not try a browser built just for discriminating people with fabulous taste like yourself? omniweb5β¦, manufacturer - NaN, price - 23.99
@@
I.3 Product Relation Prediction ()
@@ I.4 Sentiment Analysis ()
Carefully assess the user's review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
@@
This visor CD holder is a great addition for any car. It folds shut so CDs don't slide out. It fits well on any visor.
@@ I.5 Sequential Recommendation ()
1st: Bbox A392-10CP Dual 10" Sealed Carpeted Subwoofer Enclosure - Fits 1999-2007 Ford F250/350/450 Crew Cab. Electronics. Car & Vehicle Electronics. b.box.
@@
Q: GreenLife Soft Grip 11" Ceramic Non-Stick Open Wok, Burgundy. Home & Kitchen. Kitchen & Dining. GreenLife.
@@
R: Bulova Frank Lloyd Wright Luxfer Prism Wall Clock, 14", Bronze. Home & Kitchen. Home Dcor. Bulova.
@@
I.6 Multiclass Product Classification ()
@@
I.7 Product Substitute Identification ()
@@
I.8 Query Product Ranking ()
@@ I.9 Answerability Prediction ()
Document 1: I purchased two of these locks and within two uses with each one I had trouble opening the lock. The third time and I cannot get either opened at all. They are stuck on the tree. I'll order a different brand and then go cut them off. They keys each turn but they won't come off. Wish I had read the reviews.
@@
Document 3: I like the Python. It's easy to use & I feel it is well made. The problem is I use multiple game cameras & you can not buy the locks keyed alike. My local locksmith who is a Master Lock dealer said can not order them alike. I called Master Locks corporate office-and was told the same. If you need one they're great. I'm forced to look elsewhere.
@@ I.10 Answer Generation ()
Document 1: Not going to write a full review here, it's a cheap motherboard, it's $50, expect to get what you pay for. I bought this from another retailer, but wanted everyone to know that this board does in-fact support 16GB of RAM. I am using this board in a home lab VMware ESXi host, and I needed 16GB at least. Here is what you need to do: (1) Update bios to at least 1303 (here is linky) [β¦] (2) Use RAM on supported RAM list (here is linky)[β¦] I am using 16GB of Corsair VengeanceCorsair Vengeance 16GB (2x8GB) DDR3 1600 MHz (PC3 12800) Desktop Memory (CMZ16GX3M2A1600C10)and it works like a champ.
Document 2: I bought this mobo to replace my ASUS MVA2 HDMI series when upgrading my system. As said in other reviews it can definitely use a 125 watt CPU as I checked the manual before I started my upgrade. Originally this was designed for Phenom 2 X6 cpus which run over 95 watts typically on load.On my rig I'm using an AMD FX8350, 8gb of Kingston DDR31600 ram (2 sticks of 4gb in each slot dual channel), ASUS AMD Radeon 7750 1gb, Kingston 120gb ssd, 1tb Western Digital hdd, Windows XP 32bit and 7 64 bit, in an Antec desktop case with an upgraded Antec 660 watt power supply.
Document 3: This is the PLUS version and will take a 125W processor (I have a AMD 965 quad core Deneb). It will take 8GB RAM in each of the two slots. It's a good board with all the Asus extras on it. Some considerations - the legacy connectors on the back might be useful for you or not. Most people would prefer extra USB's (or a couple of fast USB). The front panel connector is on the side next to the USB connectors so if you have a short panel connector cord(s) that *might* be an issue for you.
Document 4: Always check the manufacturer's website for reliable specs. I built a computer using this board and it works perfectly.Reviewer who called this mobo ``junk" and gave it a ``1" was attempting to use a CPU this board isn't rated for- board rated for up to 95W cpu. No mobo will run with wrong RAM, CPU, inadequate power supply, etc. Double check all components, OS, peripherals, BIOS version and settings, etc. for compatibility.Make sure you know what you're doing before attempting to build your own computer or you can end up wasting a lot of money.
@@
Ground truth: You can use 8Gb modules, there's only 2 DIMMS. I'm using Corsair DDR 3 16GB kit of 2 8Gb1600Mhz

### v2

@@ Appendix H OOD Evaluation
H.1 Results on Unseen Instructions
TableΒ A17 exhibits the results of evaluating on OOD test sets with unseen instruction. As shown in tableΒ A17, models perform better when training on with diverse instructions. fine-tuned on with diverse instruction enhances the generalizability of models to the unseen instruction.
@@ H.1 Results on Unseen Instructions
In this table, βsingleβ and βdiverseβ indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model tuned over single and diverse instructions on each task is in bold.
H.2 Results on Different Base Models
TableΒ A18 shows the OOD evaluation results of with different base models. With high-quality dataset, models achieve good performance under different base models.
@@ H.2 Results on Different Base Models
In this table, βBase Modelβ presents the base models used in models. On each task, the best performance of , , and when using different base models is in bold.
H.3 Results on Models Tuned using All Tasks and Individual Tasks
As demonstrated in TableΒ A19, when generalizing to OOD test, fine-tuned on every single task can better transfer the knowledge to the specific task than fine-tuned on the task combination.
| Model | Training Tasks | ||||||
@@ H.3 Results on Models Tuned using All Tasks and Individual Tasks
In this table, βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold.
Appendix I Complete Results for the Analysis of Training Data Size
@@
This section exhibits the comprehensive results for analyzing how the training data size influences the performance of models. With the total 92K samples in , we assess data scaling with sizes of 1K, 10K, and 47K. The 1K and 10K samples are collected by randomly selecting 0.1K and 1K samples, respectively, from the training set of each of the 10 tasks. The 47K samples are assembled by randomly selecting 5K samples from each task except for , for which all its 2K training samples are included. As shown in TableΒ A20 and TableΒ A21, performances exhibit an upward trend along with the increase in data size. Notably, the F1 score of the task experiences an initial drop followed by a subsequent rise with the data size increasing. This behavior could be attributed to the imbalanced data, making the models tend to randomly guess when the data size is small. As the data size increases, the models would predict the dominant label, leading to a temporarily low F1 score. With continued data growth, the models acquire sufficient knowledge to generate accurate predictions. In general, the large-scale training data plays a crucial role in developing effective e-commerce LLMs, underscoring the significance of our extensive, comprehensive, and high-quality e-commerce instruction dataset .
Appendix J Case Study
We show the predictions from and the best-performing general-purpose LLM GPT-4 Turbo on each of the 10 tasks. From the following examples, we note that by instruction tuning using our high-quality , could outperform GPT-4 Turbo in distinguishing products (e.g. ), estimating usersβ preferences (e.g. ), and understanding products (e.g. ). For example, as shown in SectionΒ J.2, could accurately predict whether two products are the same given their title, description, manufacturer, and price, while GPT-4 Turbo struggles. In addition, the example in SectionΒ J.5 indicates that compared to GPT-4 Turbo, could better estimate usersβ preferences based on their historical interactions. Moreover, the example in SectionΒ J.6 shows that could correctly understand that DJ light is equivalent to light that will change based on the music. In contrast, without instruction tuning, GPT-4 Turbo may lack this knowledge and fail the example.
J.1 Attribute Value Extraction ()
@@
Product 2: title - omni web 5.0, description - sure you can use a standard web browser with standard features. but you didnβt choose a standard software experience - you chose the mac. why not try a browser built just for discriminating people with fabulous taste like yourself? omniweb5β¦, manufacturer - NaN, price - 23.99
@@
J.3 Product Relation Prediction ()
@@
Carefully assess the userβs review for any strong expressions of sentiment, either positive or negative. Based on your analysis, select the most fitting sentiment option from the provided list as output.
@@
This visor CD holder is a great addition for any car. It folds shut so CDs donβt slide out. It fits well on any visor.
@@
1st: Bbox A392-10CP Dual 10β Sealed Carpeted Subwoofer Enclosure - Fits 1999-2007 Ford F250/350/450 Crew Cab. Electronics. Car & Vehicle Electronics. b.box.
@@
Q: GreenLife Soft Grip 11β Ceramic Non-Stick Open Wok, Burgundy. Home & Kitchen. Kitchen & Dining. GreenLife.
@@
R: Bulova Frank Lloyd Wright Luxfer Prism Wall Clock, 14β, Bronze. Home & Kitchen. Home Dcor. Bulova.
@@
J.6 Multiclass Product Classification ()
@@
J.7 Product Substitute Identification ()
@@
J.8 Query Product Ranking ()
@@
Document 1: I purchased two of these locks and within two uses with each one I had trouble opening the lock. The third time and I cannot get either opened at all. They are stuck on the tree. Iβll order a different brand and then go cut them off. They keys each turn but they wonβt come off. Wish I had read the reviews.
@@
Document 3: I like the Python. Itβs easy to use & I feel it is well made. The problem is I use multiple game cameras & you can not buy the locks keyed alike. My local locksmith who is a Master Lock dealer said can not order them alike. I called Master Locks corporate office-and was told the same. If you need one theyβre great. Iβm forced to look elsewhere.
@@
Document 1: Not going to write a full review here, itβs a cheap motherboard, itβs $50, expect to get what you pay for. I bought this from another retailer, but wanted everyone to know that this board does in-fact support 16GB of RAM. I am using this board in a home lab VMware ESXi host, and I needed 16GB at least. Here is what you need to do: (1) Update bios to at least 1303 (here is linky) [β¦] (2) Use RAM on supported RAM list (here is linky)[β¦] I am using 16GB of Corsair VengeanceCorsair Vengeance 16GB (2x8GB) DDR3 1600 MHz (PC3 12800) Desktop Memory (CMZ16GX3M2A1600C10)and it works like a champ.
Document 2: I bought this mobo to replace my ASUS MVA2 HDMI series when upgrading my system. As said in other reviews it can definitely use a 125 watt CPU as I checked the manual before I started my upgrade. Originally this was designed for Phenom 2 X6 cpus which run over 95 watts typically on load.On my rig Iβm using an AMD FX8350, 8gb of Kingston DDR31600 ram (2 sticks of 4gb in each slot dual channel), ASUS AMD Radeon 7750 1gb, Kingston 120gb ssd, 1tb Western Digital hdd, Windows XP 32bit and 7 64 bit, in an Antec desktop case with an upgraded Antec 660 watt power supply.
Document 3: This is the PLUS version and will take a 125W processor (I have a AMD 965 quad core Deneb). It will take 8GB RAM in each of the two slots. Itβs a good board with all the Asus extras on it. Some considerations - the legacy connectors on the back might be useful for you or not. Most people would prefer extra USBβs (or a couple of fast USB). The front panel connector is on the side next to the USB connectors so if you have a short panel connector cord(s) that *might* be an issue for you.
Document 4: Always check the manufacturerβs website for reliable specs. I built a computer using this board and it works perfectly.Reviewer who called this mobo βjunkβ and gave it a β1β was attempting to use a CPU this board isnβt rated for- board rated for up to 95W cpu. No mobo will run with wrong RAM, CPU, inadequate power supply, etc. Double check all components, OS, peripherals, BIOS version and settings, etc. for compatibility.Make sure you know what youβre doing before attempting to build your own computer or you can end up wasting a lot of money.
@@
Ground truth: You can use 8Gb modules, thereβs only 2 DIMMS. Iβm using Corsair DDR 3 16GB kit of 2 8Gb1600Mhz